# Evaluating Conversational Recommender Systems with Large Language Models

### v1

@@ Abstract.
copyright: noneconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
In recent years, conversational recommendation have gained growing attention within the research community of recommender systems, and various implementations of Conversational Recommender Systems (CRSs) have been proposed (e.g.,  (Chen et al., 2019; Wang et al., 2022a; Sun and Zhang, 2018; Tu et al., 2022; Wang et al., 2022b)). Unlike traditional Recommender Systems (RS), which model user interests and generate recommendations based on implicit feedback signals such as clicking or browsing history  (Schafer et al., 2007; Ekstrand et al., 2011; Koren, 2008; Bell and Koren, 2007; Koren et al., 2009; Covington et al., 2016), CRSs solicit and identify user interests through natural language conversations to make recommendations (Zamani et al., 2023; Jannach et al., 2021; Gao et al., 2021).
In conversational recommendation, a typical conversation between a user and a CRS is shown in Figure 1. The system generates text responses to the user and also provides an idependent list of recommended items. Different from traditional RS which are solely concerned with item recommendation, CRS not only address item recommendation but also manage conversations, i.e., engaging in multi-turn, semantically coherent dialogues.
@@
In recent years, as large language models (LLMs) have demonstrated significant natural language understanding capabilities, researchers have proposed various methods to leverage LLMs for evaluating text quality and have demonstrated that LLMs have the potential to align with human preferences. (Liu et al., 2023; Fu et al., 2024; Chen et al., 2023b; Chiang and Lee, 2023; Wang et al., 2024; Gao et al., 2023). However, to our knowledge, the use of LLMs for evaluating CRS remains quite limited.
To advance the development of user-centric conversational recommendation systems, in this work, we build upon existing literature and propose a framework of evaluation metrics for CRSs using LLMs from a user-centric perspective, considering four dimensions: (1) Dialogue Behavior, which primarily focuses on whether the CRS’s actions during interactions with users meet user expectations, whether it can correct errors based on user requests, and whether it can proactively inquire about user preferences. (2) Language Expression, which primarily focuses on whether the text generated by the CRS conforms to natural human language usage, whether there are any obvious grammatical errors, and whether the language used is appropriate. (3) Recommendation Items, which primarily focuses on whether the items recommended by the recommendation system meet the user’s needs and whether they also balance diversity and novelty based on this foundation. (4) Response Content, which primarily focuses on whether the text generated by the CRS during the dialogue is semantically closely related to the recommendation list provided by the CRS, whether it offers in-depth explanations of the recommendations within the text, and whether the content of the text is factually correct. Figure 1 presents the four evaluation dimensions we proposed and the corresponding metrics under each dimension.
@@
We evaluated BARCOR (Wang et al., 2022a), CHATCRS (Wang et al., 2023a), KBRD (Chen et al., 2019), and UniCRS (Wang et al., 2022b) across the 12 factors in the proposed framework, using three LLMs: GPT-4o-mini-0718, LLaMa-3-8B, and GLM-4-Air, on the OpenDialKG (Moon et al., 2019) and ReDial (Li et al., 2019) datasets. Overall, we found that: The three evaluators generally give CHATCRS the highest score, BARCOR the second-highest score, while the rankings for KBRD and UniCRS vary depending on the factors and datasets.
@@
To our knowledge, we believe we are the first to propose a user-centric LLM-based evaluating framework for CRSs, incorporating insights from previous work in the field of recommender system, psychology, human-computer interaction. The main contributions of our work is that we systematically compiles and summarizes a comprehensive and diverse set of metrics for evaluating CRSs from a user-centric perspective, utilizing LLM-based standardized assessments. By using LLMs as agents for real users in multi-facet evaluations, our framework can consider various factors that influence user satisfaction and score CRS performance from different perspectives, bridging the gap between the fields of human-computer interaction and recommender system technology.
@@ 2. Related Work
2.1. Conversational Recommender Systems
Conversational recommender systems (CRSs) interact with users through multiple turns of dialogue, solicit and identify user interests through natural language conversations (Jannach et al., 2021; Gao et al., 2021; Zamani et al., 2023). Most of the implementation methods are based on neural networks and combine techniques such as knowledge graphs (Chen et al., 2019; Zhou et al., 2020; Wang et al., 2022a, b). In recent years, methods that directly utilize Large Language Models (LLMs) as CRS have also emerged (Wang et al., 2023a).
In these implementations, the interaction patterns between users and CRSs can be divided into two main categories: One line of work relies on handcrafted response templates and predefined interactive actions (e.g., asking users’ preferences for item attributes or making recommendations) to interact with users (e.g., (Lei et al., 2020; Tu et al., 2022; Zhou et al., 2020; Shang et al., 2023)); Another line of work focuses on interacting with users through more flexible natural language conversations, also known as chit-chat (e.g., (Chen et al., 2019; Wang et al., 2022b, a)). In our study, we primarily focus on the evaluation of chit-chat CRS.
We selected four representative CRSs for evaluation, The three methods are based on neural networks, namely BARCOR (Wang et al., 2022a), KBRD (Chen et al., 2019), and UniCRS (Wang et al., 2022b), while one directly utilizes GPT as a CRS (CHATCRS) (Wang et al., 2023a). We selected these systems because many studies ( e.g., (Wang et al., 2023a, b; Zhu et al., 2024b, a)) have used them as baselines or evaluated them, which helps to connect our work with existing evaluations of conversational recommender systems.
2.2. Evaluating Conversational Recommender Systems
In fields such as Human-Computer Interaction (HCI) and psychology, there have been many research based on user studies analyzing what factors can influence users’ experience with CRS (Jannach, 2022; Cai et al., 2022; Jin et al., 2021; Cai and Chen, 2020). Nevertheless, only a few work have proposed metrics for evaluating CRSs (Yang et al., 2024; Huang et al., 2024).
In the RS community, most work still rely on metrics such as Recall and nDCG to measure the effectiveness of CRS recommendations, and metrics like BLEU and distinct-n to measure the quality of text generated by CRS.
Since obtaining interaction logs and feedback from real users is costly, a common method for evaluating Chit-chat CRS is user simulation (Wang et al., 2023a; Zhang and Balog, 2020). Wang et al. (Wang et al., 2023a) proposed a user simulator based on LLM and argued that it can be a reliable alternative to the human evaluator. In this work, we also adopt the user simulation-based approach for evaluation, using LLMs as user simulators to engage in multi-turn dialogues with the CRS, following the work of Wang et al. (Wang et al., 2023a).
3. CoRE: Our Proposed Method
@@
We use LLMs as evaluators, providing scores from 0 to 4 based on the 12 factors mentioned above. Figure 2 illustrates the overall process of our method; Figure 1 shows how the 12 factors correspond to elements in the user-CRS interaction.
@@ 4.1. Experimental Setup
Dataset. Following previous work (e.g., (Wang et al., 2023a, b; Zhu et al., 2024b, a)), we select ReDial (Li et al., 2019) and OpenDialKG (Moon et al., 2019) as the benchmark datasets. ReDial primarily focuses on movie recommendations, and OpenDialKG encompasses a broader range of domains, including movies, books, sports, and music. Table 1 shows a summary of the two datasets.
CRS Model. Following previous work (e.g., (Wang et al., 2023a, b; Zhu et al., 2024b, a)), we present a comparative evaluation and analysis towards four CRS: BARCOR (Wang et al., 2022a), KBRD (Chen et al., 2019), and UniCRS (Wang et al., 2022b) and CHATCRS (Wang et al., 2023a). For BARCOR, KBRD, and UniCRS, we use the pre-trained models provided by Wang et al (Wang et al., 2023a); For CHATCRS, we implemented several modifications, utilizing the GPT-4o-mini-0718 as the LLM model.
User Simulator. Building on the work of Wang et al. (Wang et al., 2023a), we use GPT-4o-mini-0718 to simulate users interacting with CRS in dialogues. To ensure sufficient system dialogue text, we modified the interaction rules and prompts to ensure that each dialogue session between the user simulator and CRS includes at least three and no more than five rounds of interaction. More specifically, before each dialogue session begins, we input the ground truth item from a real user into the simulator, requiring the simulator to describe relevant attributes in the dialogue without directly revealing the ground truth item. In each round of dialogue, the system generates response text and a list of recommended items. If the system identifies the ground-truth item, and more than three rounds of dialogue have occurred, the session ends; if more than five rounds of dialogue have occurred and the system has not identified the ground-truth item in any of the previous rounds, the session also ends.
@@ 4.2. Experimental Result
Table 23, and  4 shows the evaluation results of four systems on two datasets are presented separately, using GPT-4o-mini-0718, LLaMa3-8B, and GLM-4-Air as evaluators.
@@
From the perspective of the datasets, we found that GPT and LLaMa tend to give more consistent rankings on OpenDialKG, while GPT and GLM tend to give more consistent rankings on ReDial.
Overall, we found that: The three evaluators generally give CHATCRS the highest score, BARCOR the second-highest score, while the rankings for KBRD and UniCRS vary depending on the factors and datasets. We also found that: In most cases, the score rankings given by GPT and LLaMa are consistent; but the score rankings given by GLM are not as consistent with those of GPT or LLaMa.
@@ 5.1. Main Findings
Overall, we found that: The three evaluators generally give CHATCRS the highest score, BARCOR the second-highest score, while the rankings for KBRD and UniCRS vary depending on the factors and datasets. We also found that: In most cases, the score rankings given by GPT and LLaMa are consistent; but the score rankings given by GLM are not as consistent with those of GPT or LLaMa.
We believe there are two possible reasons for the inconsistency in LLMs’ scores. First, each model learns different parameterized knowledge during training due to variations in the training corpora, which can affect their judgment. Second, some factors, such as diversity or novelty, are more subjective in their criteria compared to others, such as groundness or grammar, which can also influence the scores given by LLMs, reflected in the system rankings and the standard deviation of each system’s score.
@@ 5.2. Limitations and Future Directions
Firstly, our work has not examined how well LLM evaluators’ scores align with human ratings, and future research should include user studies and human feedback. Secondly, we have not yet addressed the gap in CRS evaluation where recommendation and conversational tasks are treated separately and need to be integrated into a unified approach.
Additionally, previous work on search system evaluation has shown that users may be influenced by cognitive biases during their interactions with systems (Chen et al., 2023a, 2024). Future research needs to further integrate psychological theories to make the evaluation results closer to real user feedback.
@@ A.1. Instruction Template
Figure 3 shows the instruction templates we provided to the LLM evaluators. The content of FACTOR_DEFINITION, EVALUATION_STANDARD, EVALUATION_STEP of each factor and an example of CONVERSATION_LOG can be found below.
@@ A.2. Evaluation Guidelines
Figure 4567,89101112131415 shows the evaluation guidelines of each factor we provided to the LLM evaluators respectively.

### v2

# Large Language Models as Evaluators for Conversational Recommender Systems

@@ Abstract.
Conversational Recommender Systems, Information Access, Dialogue Systems,Information System Evaluation
copyright: noneconference: The 1st Workshop on Human-Centered Recommender Systems (HCRS) at TheWebConf 2025; 28 April - 2 May, 2025; Sydney, Australiaisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
@@ 1. Introduction
In recent years, conversational recommendation have gained growing attention within the research community of recommender systems, and various implementations of Conversational Recommender Systems (CRSs) have been proposed (Chen et al., 2019; Wang et al., 2022a; Sun and Zhang, 2018; Tu et al., 2022; Wang et al., 2022b). Unlike traditional Recommender Systems (RS), which model user interests and generate recommendations based on implicit feedback signals such as clicking or browsing history  (Koren, 2008; Mao et al., 2021; Zhu et al., 2023, 2022), CRSs solicit and identify user interests through natural language conversations to make recommendations (Zamani et al., 2023; Jannach et al., 2021; Gao et al., 2021).
@@
In conversational recommendation, a typical conversation between a user and a CRS is shown in Figure 1. The system generates text responses to the user and also provides an idependent list of recommended items. Different from traditional RS that focus solely on item recommendation, CRS not only provide item recommendation but also manage conversations, i.e., engaging in multi-turn, semantically coherent dialogues.
@@
In recent years, large language models (LLMs) have shown significant capabilities in natural language understanding. As a result, researchers have proposed various methods for leveraging LLMs to evaluate text quality, demonstrating their potential to align closely with human preferences (Liu et al., 2023; Fu et al., 2024; Chen et al., 2023b; Chiang and Lee, 2023; Wang et al., 2024; Gao et al., 2023). However, to the best our knowledge, the application of LLMs for evaluating CRS remains relatively underexplored.
To advance the development of conversational recommendation systems, this work builds upon existing literature and proposes a user-centric evaluation framework based on LLMs, which can comprehensively and naturally quantify users’ satisfaction with the interacted CRS. Specifically, within the framework, we first generate interaction dialogues between users and the system by substituting real humans with user simulators to avoid high cost as (Wang et al., 2023a). We then evaluate the proposed metrics based on the logs by prompting the LLM.
Our evaluation framework consists of 12 evaluation metrics of four dimensions, including dialogue behavior, language expression, recommendation items, and response content, as shown in Figure 1. We describe these four dimensions as follows:
  • Dialogue Behavior. This dimension focuses on whether the CRS’s actions during interactions align with user expectations, its ability to correct errors based on user requests, and whether it proactively inquires about user preferences.
  • Language Expression. This dimension evaluates whether the text generated by the CRS follows natural human language usage, is free of obvious grammatical errors, and whether the language is appropriate in context.
  • Recommendation Items. This dimension assesses whether the items recommended by the CRS meet the user’s preferences and needs while balancing diversity and novelty.
  • Response Content. This dimension examines whether the text generated by the CRS during the dialogue is semantically relevant to the recommendation list, whether it provides in-depth explanations of the recommendations, and whether the content is factually accurate.
@@
We conducted evaluations of BARCOR (Wang et al., 2022a), CHATCRS (Wang et al., 2023a), KBRD (Chen et al., 2019), and UniCRS (Wang et al., 2022b) according to the 12 metrics within the proposed framework. The evaluations were carried out using three LLMs, namely GPT-4o-mini-0718, LLaMa-3-8B, and GLM-4-Air, on the OpenDialKG (Moon et al., 2019) and ReDial (Li et al., 2019) datasets. Overall, our findings indicate that: generally, the three evaluators assign the highest score to CHATCRS, followed by BARCOR with the second highest score. However, the rankings of KBRD and UniCRS fluctuate based on the specific factors and datasets under consideration.
@@
To the best of our knowledge, we are the first to propose a user-centric evaluation framework for CRSs based on LLMs. This framework incorporates valuable insights from previous research in the fields of recommender systems, psychology, and human-computer interaction. The main contributions of our work are as follows: First, we have systematically compiled and summarized a comprehensive and diverse set of metrics for evaluating CRSs from a user-centric perspective. These metrics are evaluated through LLM-based standardized assessments. Second, by employing LLM-based user simualtors as surrogates for real users in multi-faceted evaluations, our framework can take into account a wide range of factors that influence user satisfaction in a highly efficient manner. It is thus able to score the performance of CRSs from different viewpoints, effectively bridging the gap between the fields of human-computer interaction and recommender systems.
@@ 2. Related Work
2.1. Conversational Recommender Systems
Conversational recommender systems (CRSs) interact with users through multiple turns of dialogue, soliciting and identifying user interests through natural language conversations (Jannach et al., 2021; Gao et al., 2021; Zamani et al., 2023). Most implementation methods are based on neural networks and incorporate techniques such as knowledge graphs (Chen et al., 2019; Zhou et al., 2020b; Wang et al., 2022a, b). In recent years, methods that directly utilize Large Language Models as CRSs have also emerged (Wang et al., 2023a).
In these implementations, the interaction patterns between users and CRSs can be classified into two main categories. One approach relies on handcrafted response templates and predefined interactive actions (e.g., asking users about item attributes or making recommendations) to engage with users (Lei et al., 2020; Tu et al., 2022; Zhou et al., 2020b; Shang et al., 2023). The other approach focuses on more flexible interactions through natural language conversations, commonly referred to as chit-chat (Chen et al., 2019; Wang et al., 2022b, a). In our study, we primarily focus on evaluating chit-chat CRSs.
We selected four representative CRSs for evaluation. Three of these methods are based on neural networks, i.e., BARCOR (Wang et al., 2022a), KBRD (Chen et al., 2019), and UniCRS (Wang et al., 2022b), while one directly utilizes GPT as a CRS, namely CHATCRS (Wang et al., 2023a). These systems were chosen because they have been widely used as baselines or evaluated in previous studies (e.g., (Wang et al., 2023a, b; Zhu et al., 2024b, a)), which allows our work to build upon and connect with existing evaluations of conversational recommender systems.
2.2. Evaluating Conversational Recommender Systems
In fields such as Human-Computer Interaction (HCI) and psychology, numerous studies have analyzed the factors influencing users’ experiences with CRSs (Jannach, 2022; Cai et al., 2022; Jin et al., 2021; Cai and Chen, 2020). However, few studies have formally proposed metrics for evaluating CRSs (Yang et al., 2024; Huang et al., 2024).
In the community of recommender systems, traditional evaluation methods for CRSs typically focus on assessing recommendation accuracy (with metrics such as Recall) and nDCG and dialogue response quality (with metrics such as BLEU and distinct-n) in isolation (Wang et al., 2022a; Zhou et al., 2020a). This approach significantly overlooks the role of the dynamic interaction process in shaping user experiences, thereby failing to accurately capture the true performance of a CRS (Wang et al., 2023a). Although online user testing is considered the gold standard for evaluating CRSs, due to its ability to capture real user interactions and feedback, its high time demands and substantial labor costs make it difficult to scale (Gao et al., 2021).
To tackle this problem, user simulators have emerged as a substitute of real humans to interact with CRSs for performance evaluation (Zhang and Balog, 2020; Afzali et al., 2023; Chen et al., 2025). Recently, LLM-based user simulators have been continuously proposed, including single-prompt methods (Wang et al., 2023a; Zhu et al., 2024b; Kim et al., 2024) and agent-based architectures (Zhu et al., 2024a; Chen et al., 2025). However, these existing simulator-based methods only use a very limited set of metrics, which is insufficient to comprehensively evaluate a CRS. In this work, we build upon these advancements by leveraging LLM-based user simulators in our evaluation framework, and further propose a comprehensive evaluation metrics to evaluate the interaction performance based on user-system interaction dialogues. We also propose to use LLMs as evluators for these metrics.
3. The Proposed Method
In this paper, we propose a user-centric evaluation framework based LLMs for conversational recommender systems as shown by Figure 2. It first employs LLM-based user simulators to interact with CRS under various contexts to obtain a diverse set of interaction dialogues, and then utilizes LLM as evaluator to score the carefully designed evaluation metrics (ranging from 0 to 4 in our experiments).
@@ 3. The Proposed Method
We use LLMs as evaluators to assign scores from 0 to 4 for the 12 factors mentioned above separately, based on the carefully designed prompts as detailedly presented in Appendix A.
@@ 4.1. Experimental Setup
Dataset. Following previous work (Wang et al., 2023a, b; Zhu et al., 2024b, a), we select ReDial (Li et al., 2019) and OpenDialKG (Moon et al., 2019) as the benchmark datasets. ReDial primarily focuses on movie recommendations, and OpenDialKG encompasses a broader range of domains, including movies, books, sports, and music. Table 1 shows a summary of the two datasets.
CRS Model. Following previous work (Wang et al., 2023a, b; Zhu et al., 2024b, a), we present a comparative evaluation and analysis towards four CRSs, including BARCOR (Wang et al., 2022a), KBRD (Chen et al., 2019), UniCRS (Wang et al., 2022b) and CHATCRS (Wang et al., 2023a). For BARCOR, KBRD, and UniCRS, we use the pre-trained models provided by Wang et al (Wang et al., 2023a). For CHATCRS, we implemented several modifications with the GPT-4o-mini-0718 as the LLM model.
User Simulator. Building on the work of Wang et al. (Wang et al., 2023a), we use GPT-4o-mini-0718 to simulate user interactions with the CRS in dialogues. To ensure sufficient user-system interactions, we modified the interaction rules and prompts to guarantee that each dialogue session between the user simulator and CRS includes at least three and no more than five rounds of interaction. More specifically, before each session begins, we input the ground-truth item from a real user into the simulator, instructing it to describe relevant attributes during the dialogue without directly revealing the ground-truth item. In each round of dialogue, the system generates a response and a list of recommended items. The session ends if the system identifies the ground-truth item and more than three rounds of dialogue have occurred. Alternatively, the session also ends if more than five rounds of dialogue have occurred and the system has not identified the ground-truth item in any of the previous rounds.
@@ 4.2. Experimental Result
Table 23, and 4 show the evaluation results of four systems on two benchmark datasets, using GPT-4o-mini-0718, LLaMa3-8B, and GLM-4-Air as evaluators.
@@
From the perspective of the datasets, we observed that GPT and LLaMa tend to provide more consistent rankings on OpenDialKG, while GPT and GLM show greater consistency on ReDial.
@@ 5. Conclusion
In this paper, we have proposed a user-centric evaluation framework based on LLMs for accurately and comprehensively evaluating conversational recommender systems. In this framework, we leverage LLM-based user simulators to interact with CRSs to obtain the interaction dialogues, and then used LLM as evaluators to assign scores to the 12 evaluation metrics of four dimensions based on the interaction dialogues. We further conduct experiments with the framework on four CRSs and two benchmark datasets to demonstrate its performance.
@@ 5.1. Main Findings
We found that the three evaluators generally awarded the highest score to CHATCRS, the second-highest to BARCOR, while the rankings for KBRD and UniCRS varied depending on the factors and datasets. Additionally, we found that in most cases, the score rankings from GPT and LLaMa were consistent; however, the rankings from GLM were less consistent with those of GPT and LLaMa.
We believe there are two possible reasons for the inconsistencies in the LLMs’ scores. First, each model learns different parameterized knowledge during training due to variations in the training corpora, which can affect its judgments. Second, certain factors, such as diversity and novelty, are more subjective in nature compared to others, such as groundedness or grammar. Such subjectivity can also influence the scores given by the LLMs, as reflected in the system rankings and the standard deviation of each system’s scores.
@@ 5.2. Limitations and Future Directions
As a pioneering work in evaluating conversational recommendation systems based on Large Language Models, this study still has several limitations that need to be addressed. Firstly, we have not examined the alignment between LLM evaluators’ scores and human ratings. Future research should incorporate user studies and human feedback to bridge this gap. Secondly, we have not yet addressed the gap in CRS evaluation where recommendation and conversational tasks are treated separately. These tasks need to be integrated into a unified approach. Additionally, previous work on search system evaluation has shown that users may be influenced by cognitive biases during their interactions with systems (Chen et al., 2023a, 2024). Future research needs to further integrate psychological theories to make the evaluation results closer to real user feedback.
@@ A.1. Instruction Template
Figure 3 shows the instruction templates provided to the LLM evaluators. The content of FACTOR_DEFINITION, EVALUATION_STANDARD, EVALUATION_STEP for each factor, and an example of CONVERSATION_LOG can be found below.
@@ A.2. Evaluation Guidelines
Figure 4567891011121314, and 15 show the evaluation guidelines of each factor we provided to the LLM evaluators, respectively.